### v1

# Defending Against Malicious Behaviors in Federated Learning with Blockchain

Abstract
In the era of deep learning, federated learning (FL) presents a promising approach that allows multi-institutional data owners, or clients, to collaboratively train machine learning models without compromising data privacy. However, most existing FL approaches rely on a centralized server for global model aggregation, leading to a single point of failure. This makes the system vulnerable to malicious attacks when dealing with dishonest clients. In this work, we address this problem by proposing a secure and reliable FL system based on blockchain and distributed ledger technology. Our system incorporates a peer-to-peer voting mechanism and a reward-and-slash mechanism, which are powered by on-chain smart contracts, to detect and deter malicious behaviors. Both theoretical and empirical analyses are presented to demonstrate the effectiveness of the proposed approach, showing that our framework is robust against malicious client-side behaviors.
{IEEEImpStatement}
Federated learning has been a promising solution to utilize multi-site data while preserving users’ privacy. Despite the success of integrating blockchain with federated learning to decentralize global model aggregation, the protection of this integration from clients with malicious intent in federated scenarios remains unclear. This paper presents the first formulation of this problem and the proposed stake-based aggregation mechanism shows robustness in detecting malicious behaviors. The results in this work not only pose a new research direction in federated learning, but can also benefit a wide variety of applications such as finance and healthcare.
{IEEEkeywords}
Blockchain, Deep Learning, Federated Learning, Trustworthy Machine Learning
1 Introduction
[…]
Nowadays, machine learning (ML), or more specifically, deep learning, has transformed a broad spectrum of industries, ranging from finance to healthcare. In current ML paradigms, training data are first collected and curated, and then ML models are optimized by minimizing certain loss criteria on the training data. A common underlying assumption in the learning environment is that the training data can be instantly accessed or easily distributed across computing nodes without communication constraints, i.e. missing data are centralized.
However, in a system with multiple clients (i.e. missing data holders), to ensure data centralization, clients have to upload local data to a centralized device (e.g. missing a central server) to conduct the centralized training described above. Despite the success of centralized training in various deep learning applications [1, 2, 3], there is growing concern about data privacy and security, especially when the local data held by the clients are private or contain sensitive information. Especially, to ensure data governance, strict data regulations have been established [4, 5].
Figure 1: A stake-based aggregation mechanism for FL with blockchain. In each round, the proposers are randomly selected from the participating clients to perform local training and upload local updates to the blockchain. Then, voters download the aggregated local updates from the blockchain, perform local validation, and vote for acceptance or rejection. If the majority of voters vote for accepting the global aggregation, the global model will be updated, and the proposers and the voters who vote for acceptance will be rewarded. Conversely, if the majority of voters vote for rejection, the global model will not be updated, and the proposers and the voters who vote for acceptance will be slashed.
To address the aforementioned concern, federated learning (FL) has been proposed [6]. In a typical FL system, a central server [7] is responsible for aggregating and synchronizing model weights, while a set of clients manipulate multi-site data. This facilitates data governance, as clients only exchange model weights or gradients with a central server instead of uploading local data to the central server, and has led to FL becoming a standardized solution to utilize multi-site data while preserving privacy.
Though FL perfectly implements data decentralization, a trustworthy central server is required in the system. In such a system design, the central server in fact has privileges over clients, as the central server determines the global aggregation and synchronization. If the central server is compromised or manipulated by a malicious party, the clients are vulnerable if the central server intentionally distributes problematic model updates. This can potentially increase the cost of system management and maintenance. Towards avoiding this single point of failure, many efforts have been made to decentralize the central server, and one particularly promising solution is to use a blockchain as decentralized storage [8].
Originally proposed for cryptocurrencies, a blockchain is a distributed ledger that can record the state transition information among multiple parties [9, 10], without relying on a centralized server. Blockchain technology has gained widespread attention for its potential to revolutionize a variety of industries, such as finance [9], healthcare [11], and supply chain management [12]. By leveraging the decentralized nature of the blockchain, FL can benefit from increased security, privacy, and efficiency, as well as reduced reliance on centralized servers [13]. Concretely, in FL with blockchain, each client participating in the learning process uploads their local model updates to the blockchain, where they are stored in blocks, the metadata of a blockchain system. These blocks are then used to aggregate the local model updates into a global model, which can be downloaded by the clients. The use of blockchain smart contracts [9], which are computer programs triggered by blockchain events, ensures that the global aggregation process is performed automatically and transparently, without the need for human intervention or centralized control.
Though integrating blockchain with existing FL systems can partially solve the threat to the central server, FL systems are still vulnerable to client-side malicious attacks [14]. In this work, we define malicious behaviors as actions that intentionally decrease the learning performance (e.g. missing accuracy and convergence) of the global model. The attackers can sabotage the FL systems via attacks such as data poisoning [15] or model poisoning [14]. This work focuses on defending against client-side malicious attacks.
We propose a generic framework that can integrate an FL system with a blockchain system and can defend against malicious attacks. The proposed defense mechanism is motivated by proof-of-stake (PoS) [16], a consensus mechanism in blockchain, and The Resistance [17], a role-playing board game. PoS has an incentive mechanism that encourages honest behaviors by rewarding it and punishes dishonest behaviors via slashing. The Resistance, on the other hand, has two mismatched competing parties, where the party with a larger size is denoted as the resistance force and the other party is denoted as the spies. In The Resistance, there is a voting mechanism where, in each round, each player conducts independent reasoning and votes for a player, and the player with the highest votes will be deemed as a “spy” and kicked out of the game. The goal of the resistance force is to vote out all the spies while the spies aim to impersonate the resistance force and survive until the end. Based on these two concepts, this work proposes a novel majority-voting mechanism for global aggregation where each participating client independently validates the quality of aggregated local updates and votes for acceptance of the global update. The aggregation mechanism is stake-based where participating clients stake assets [^1] In practice, the staked assets can be linked with cryptocurrency or real currency to increase the financial cost of malicious attacks. or tokens (a quantitative measurement of the asset, which can be used to indicate the trustworthiness of the client in our system) for their own actions. There are two types of actions, proposing (uploading local updates) and voting. If the majority vote is to accept the global aggregation, a proposer will be refunded with its staked tokens and a voter who votes for acceptance will not only be refunded but also be rewarded with the staked tokens from the voters who vote for rejection, and vice versa. The overall procedure of the stake-based aggregation mechanism is illustrated in Fig. 1.
We evaluate the proposed framework on a practical financial problem, namely loan default prediction. We simulate the FL and blockchain environment for the Lending Club Kaggle challenge dataset to conduct experiments in a controllable setting and to provide insights into the problem of interest. We empirically show that an FL system can maintain robust performance under malicious attacks by introducing the proposed stake-based aggregation mechanism.
The contributions of this work are summarized as follows: {IEEEenumerate}
We formulate the problem of decentralized federated learning with blockchain in the presence of malicious attacks.
We propose a stake-based aggregation mechanism for federated learning systems that can defend against malicious attacks.
We evaluate the robustness of the proposed framework in a simulated environment and provide initial empirical insights into the problem.
2 Related Work
2.1 Federated Learning
The concept of FL comes from the necessity of on-device training, where the training data have to remain on the device [6]. The clients of FL are distributed at different physical locations that are connected to the internet, which exposes a few security risks compared with distributed learning. First, as the local dataset belongs to the client, FL has to take the users’ privacy into consideration. This can be addressed by integrating privacy-preserving techniques into FL, such as differential privacy [18]. Second, FL can be manipulated via internet access, e.g. missing the central server can be compromised by a third party. Third, a new client could participate in the federated training at any time if it meets the required criteria. This means that clients with malicious intentions can also join the federated systems while meeting the initial criteria. This work focuses on the third risk as the second risk is mitigated by replacing the central server with a blockchain. Traditional FL methods can only detect and defend the malicious clients via assigning small weights to malicious clients during global aggregation based on the divergence of learned parameters [19] or using unsupervised learning methods (e.g. missing anomaly detection [20] and clustering [21]) to get rid of malicious clients. However, none of these methods tackles the second challenge and these methods do not consider the third challenge. The proposed framework utilizes the blockchain to ensure the security of global aggregation and defends the client-side malicious behaviors with a novel majority-voting mechanism.
2.2 Blockchain
Blockchains refer to distributed ledgers that operate on a global peer-to-peer (P2P) network, as exemplified by popular cryptocurrencies such as Bitcoin [22] and Ethereum [9]. One of the defining characteristics of blockchain technology is the ability for users to freely join or leave the network, without a central authority in place to ensure common agreement on the distributed ledgers. Instead, users rely on consensus protocols [16, 23], such as proof-of-work (PoW) or PoS, to achieve agreement in a distributed setting.
Figure 2: Blockchain workflow overview. The sender broadcasts the issued transaction to the P2P network, which will be confirmed by the miners. The confirmed transaction will be stored on a public blockchain and can be read by the recipient. Blockchain miners typically adopt a consensus mechanism to achieve an agreement on the state of the blockchain.
As shown in Fig. 2, in a blockchain system, a transaction typically involves a sender who wishes to transfer a digital asset, such as a cryptocurrency, to a recipient. The sender initiates the transaction by creating a digital signature that includes the transaction details and the sender’s private key, which is used to verify the sender’s identity and authorize the transfer. The transaction is then broadcasted over a P2P network to miners, who are participants in the network responsible for verifying and adding new blocks of transactions to the blockchain. Miners validate and confirm the transaction using consensus protocols, to ensure that the transaction is legitimate and not a duplicate or fraudulent transaction. Once confirmed, the transaction is added to a block, which is then cryptographically linked to the previous block using hash functions [24], forming a chain of blocks (i.e. missing, blockchain). The block is then propagated to all the participants in the network, creating a decentralized, immutable record of the transaction. Finally, the recipient can access the digital asset by using their private key to authenticate their identity and claim ownership of the asset. The use of cryptography and consensus protocols ensures the security, transparency, and decentralization of the transaction process, making blockchain technology a promising solution for a variety of applications beyond cryptocurrency, e.g., insurance [25], healthcare [11], supply chain management [12], energy [26], and Internet of Things (IoT) [27].
Another key feature of blockchain technology is the use of smart contracts [9], which are quasi-Turing-complete programs that can be executed within a virtual machine. When a transaction is initiated, a smart contract is typically used to encode the terms and conditions of the transaction, such as the amount, currency, and time of transfer. The smart contract is then stored on the blockchain network and executed automatically when the predefined conditions are met. The execution of the smart contract verifies the transaction, ensuring that it meets the agreed-upon terms and conditions, and then automatically transfers the digital asset or currency to the recipient. Smart contracts can be leveraged to build a wide range of decentralized applications (DApps), such as decentralized finance (DeFi) services [28].
2.3 Federated Learning with Blockchain
Traditional FL faces challenges [29], such as privacy and security concerns, unreliable communication, and difficulty in reaching a consensus among the parties. Blockchain, on the other hand, provides a decentralized, secure, and transparent platform for data storage and sharing. This makes the use of blockchain for FL a promising direction to potentially address privacy and security concerns by allowing parties to keep their data private while still contributing to the training process. Additionally, blockchain can provide a secure communication channel for FL participants and ensure the integrity of the FL process.
Current blockchain-based FL designs [30, 8, 31, 32] have been broadly used in diverse fields, including mobile edge computing [33], IoT [34] and distributed machine learning [35]. Despite the potential benefits of combining FL with blockchain, several challenges remain. For instance, FL systems are still vulnerable to client-side malicious attacks [14] and lack incentive-compatible mechanisms to motivate FL participants to behave honestly during the training process. Multiple reputation-based incentive mechanisms [36, 37] have recently been proposed to encourage participants and enhance model accuracy in blockchain-based FL. However, it remains unclear how to effectively utilize the blockchain infrastructure and leverage its inherent incentive mechanism (i.e., cryptocurrencies) to incentivize trustworthy FL behaviors and penalize malicious clients.
3 Problem Formulation
This section introduces the problem of interest, the definition of the malicious behaviors considered, and the underlying assumptions in this work.
3.1 Setup
There are $K>1$ clients in a federated system. Let $\mathcal{K}=\{1,2,\cdots,K\}$ denote the set of all clients. Let $\mathcal{D}_{k}$ denote the local data stored in client $k$, we have $\mathcal{D}_{k}\cap\mathcal{D}_{l}=\emptyset$ for $k\neq l$ and $k,l\in\mathcal{K}$. Each local dataset $\mathcal{D}_{k}$ can be randomly split into a training set and a test set, which are both private to client $k$. In addition to $K$ clients, a blockchain plays the role of a parameter server [7] for global aggregation. Let $f_{\theta}$ be the model of interest. In the parameter server, the parameter set $\theta_{0}^{0}$ is randomly initialized at round 00 and $K$ clients download $\theta_{0}^{0}$ from the blockchain as $K$ local copies $\{\theta_{k}^{0}\}_{k=1}^{K}$ for full synchronization. During the federated optimization phase, a set of $\mathcal{K}_{p}^{t}$ clients is randomly selected for round $t$. For each $k\in\mathcal{K}_{p}^{t}$, the client $k$ updates $\theta_{k}^{t-1}$ by training on the training set of $\mathcal{D}_{k}$ independently for a number of local epochs. Then, the blockchain aggregates updated $\{\theta_{k}^{t}\}_{k\in\mathcal{K}}$ collected from all the $K$ clients to update $\theta_{0}^{t}$. The $K$ clients then synchronize with the parameter server, i.e. missing $\theta_{k}^{t}\leftarrow\theta_{0}^{t}$. To facilitate data governance, as required in among others the medical domain [5, 4], we assume that the patient’s data (either raw data or encoded data) in a client can not be uploaded to the blockchain or other clients, i.e. missing only parameters $\{\theta_{k}\}_{k=0}^{K}$ and metadata (e.g. missing the statistics of data) [38, 39] can be exchanged between the blockchain and the clients. It is worth mentioning that this work focuses on the interactions between FL and blockchain, where blockchain computing (or mining, in a more fashionable sense) and the application of additional privacy-preserving techniques [18] are considered orthogonal research directions and thus beyond the scope of this work.
3.2 Malicious Behaviors
The definition of malicious behavior in this work is an action that intentionally decreases the global model performance. There are two types of actions for each client that interact with the federated system, i.e. missing a client can propose (i.e. missing be a proposer) and vote (i.e. missing be a voter). Proposing is to upload local model or gradient updates to the parameter server, while voting is a peer-review process to validate the “virtually” aggregated model updates. The technical details of the two actions are described in Sec. 4. There are thus two corresponding malicious behaviors. The first malicious behavior is to propose harmful local model updates and the second one is to vote dishonestly. More specifically, in the second case, a client votes for approval when it is aware that the proposed model updates are poisoned and votes for rejection when there is no evidence that indicates that the proposed model updates are poisoned. It is worth mentioning that the clients themselves might not intentionally attack the FL system as they can be compromised by attackers. For simplicity, we define the clients that have malicious behaviors as malicious clients in this work, denoted as $\mathcal{K}_{m}$. We use $\eta$ to denote the ratio of malicious clients among all clients, i.e. missing $\eta=\frac{|\mathcal{K}_{m}|}{\mathcal{K}}$, where ||$\cdot$|| ⋅ | is the cardinality of a set.
3.3 Assumptions
There are six important assumptions in this work.
{IEEEitemize}
A1: The goal of malicious behaviors is to decrease the global model performance. This is also reflected in Sec. 3.2. Under these assumptions, behaviors that are harmful to the system but do not influence the global model performance are beyond the scope of discussion in this work. An example is eavesdropping, i.e. missing cloning the model specifications.
A2: All clients are rational. This means that both honest and malicious clients expect to maximize their gain or minimize their loss while achieving their goals.
A3: Following previous studies on blockchain [13], we assume that $\eta$ is strictly smaller than $50\%$. This means there are always more honest clients than malicious clients in a federated system.
A4: There is no collusion among malicious clients. That is to say, each malicious client acts independently. In the application scenarios of this work (e.g. missing Sec. 5), there is a minimal bar for a client to participate in the system, e.g. missing basic qualifications or industry standards for clinical or financial institutes. Meanwhile, it is difficult to compromise multiple clients with independent cybersecurity systems simultaneously. Thus, we deem that it is almost impossible to launch large-scale multi-agent attacks in the application scenarios of interest. As an exploratory study, this work considers the single-agent scenario. (See Sec. 4.5 for the discussion on a multi-agent scenario.)
A5: There is no capacity constraint on the hardware, including computing, communication, and storage, allowing us to solely focus on the algorithmic side of the problem.
A6: The underlying blockchain of the FL system of interest is running securely with a consensus protocol that ensures the validity and integrity of transactions and blocks. While the security of the blockchain is crucial for the overall security of the FL system, addressing the malicious miners falls outside the scope of this study.
4 Method
Figure 3: A round-based training process. In the initial state (indexed as ①), both honest (black) and malicious (red) clients exist in an FL system. In the final state (indexed as ⑥), all malicious clients are expected to be removed from the system. To reach the final state from the initial state, multiple rounds of training are required. Here are two possible scenarios, the proposed aggregation is either approved (the upper branch) or denied (the lower branch) by the voters. In each round (within the dotted orange line), a subset of clients are randomly selected as proposers, and another subset of clients are randomly selected as voters. The proposers and voters interact with the blockchain following the order of orange arrows (from ② to ⑤).
4.1 Federated Aggregation
In this work, we illustrate the proposed framework in the context of the seminal FL method, FedAVG [6]. At the end of round $t$, the local models $\{\theta_{k}^{t}\}_{k=1}^{K}$ are uploaded and aggregated as a weighted average:
$\theta_{0}^{t}=\sum_{k=1}^{K}a_{k}\theta_{k}^{t-1},$ (1)
where $a_{k}=\frac{n_{k}}{N}$. The metadata $n_{k}=|\mathcal{D}_{k}|$ is the number of local training examples stored in client $k$ and $N=\sum_{k=1}^{K}n_{k}$ is the total number of training examples in the $K$ clients.
4.2 Local Validation
In contrast to standard FL algorithms, the aggregated global model is not recorded in a block directly. Instead, $\tilde{\theta}_{0}^{t}$, a copy of $\theta_{0}^{t}$ is downloaded by a randomly selected set of clients, denoted as voters, $\mathcal{K}_{v}^{t}$. A voter $k$ runs a local inference with $\tilde{\theta}_{0}^{t}$ on its local test set and outputs a local validation score. The local validation score $s^{t}_{k}$ is a scalar, which can be linked with common metrics of ML tasks [^2] For example, common evaluation metrics include accuracy for classification, mean Intersection over Union (mIOU) for semantic segmentation, and mean average precision (mAP) for object detection.. If $s^{t}_{k}$ is not lower than a threshold, the voter votes for accepting this aggregated model; otherwise, the voter votes against it. The threshold can be based on a validation score $s^{t-1}_{k}$ acquired in the previous round. In the training of ML tasks, the scores can be volatile due to the characteristics of the tasks. Thus, a hyperparameter $\epsilon\in(0,1)$ is introduced to control the tolerance of performance decrease in a single round. Mathematically, Voter $k$ has the following score.
$v_{k}^{t}=\begin{cases}\phantom{+}1,&s^{t}_{k}\geq(1-\epsilon)s^{t-1}_{k}\\ -1,&s^{t}_{k}<(1-\epsilon)s^{t-1}_{k}\end{cases}$ (2)
It is worth mentioning that it is almost impossible for the attackers to manipulate scores by fooling all the randomly selected voters (e.g. missing via adversarial attacks [40]). According to A5, the majority of voters are honest. It is thus difficult to attack (either via data poisoning or model poisoning) as the validation set of each client is private.
4.3 Majority Voting
The majority voting process for whether to apply the global aggregation operation at round $t$ can be described below. Here, we use a binary variable $a^{t}$ to denote the decision.
$a^{t}=\begin{cases}\phantom{+}1,&\sum_{k\in\mathcal{K}_{v}}v_{k}>0\\ -1,&\sum_{k\in\mathcal{K}_{v}}v_{k}\leq 0\end{cases}$ (3)
If $a^{t}=1$, the global aggregation will be finalized and recorded in the block; otherwise, the global aggregation will be discarded.
4.4 Asset Redistribution
As there are two independent actions, there are two parallel reward-and-slash designs for proposing and voting. For both actions, the randomly selected proposers and voters are required to stake a fixed sum of tokens before they act. If some of these actors fail to stake (they do not have enough tokens left), they lose their access to the blockchain and are removed from the FL system permanently. Proposers will be rewarded with tokens accumulated in an independent pool (if there are any tokens left in the pool) if the global aggregation is approved and lose their stakes if the global aggregation is rejected. The reward-and-slash design for the proposers is illustrated in Algorithm 1. For the voters, the majority party will not only take back their stakes but also be rewarded with the staked tokens lost by the minority party. The reward-and-slash design for the voters is illustrated in Algorithm 2. In the following section, Sec. 4.5, we demonstrate that under the proposed design and assumptions in Sec. 3.3, malicious voters have no incentive to make dishonest votes.
1:$a^{t}$: Majority voting decision at round $t$
2:$\mathcal{K}$: Set of participating clients at round $t$
3:$\mathcal{K}_{p}^{t}$: Set of proposers at round $t$
4:$M_{k}$: Asset of client $k$
5:$\gamma_{p}$: Staked tokens for proposing
6:${pool}_{p}$: Pool for proposers
7:if $a^{t}==-1$ then
8:     for $k\in\mathcal{K}_{p}^{t}$ do
9:         if $M_{k}\geq\gamma_{p}$ then
10:              $M_{k}\leftarrow M_{k}-\gamma_{p}$
11:              ${pool}_{p}\leftarrow{pool}_{p}+\gamma_{p}$
12:         else
13:              ${pool}_{p}\leftarrow{pool}_{p}+M_{k}$
14:              $M_{k}\leftarrow 0$
15:              $\mathcal{K}^{t}\leftarrow\mathcal{K}^{t}\setminus\{k\}$
16:else
17:     if ${pool}_{p}>0$ then
18:         for $k\in\mathcal{K}_{p}^{t}$ do
19:              $M_{k}\leftarrow M_{k}+\frac{{pool}_{p}}{|\mathcal{K}_{p}^{t}|}$
20:         ${pool}_{p}\leftarrow 0$
Algorithm 1 Reward-and-slash design for a set of randomly selected proposers.
1:$a^{t}$: Majority voting decision at round $t$
2:$\mathcal{K}$: Set of participating clients at round $t$
3:$\mathcal{K}_{v}^{t}$: Set of voters at round $t$
4:$\mathcal{K}_{m}^{t}$: Set of voters at round $t$ with $v_{k}^{t}==a^{t}$
5:$M_{k}$: Asset of client $k$
6:$\gamma_{v}$: Staked tokens for voting
7:${pool}_{v}$: Pool for voters
8:for $k\in\mathcal{K}_{v}^{t}\setminus\mathcal{K}_{m}^{t}$ do
9:     if $M_{k}\geq\gamma_{v}$ then
10:         $M_{k}\leftarrow M_{k}-\gamma_{v}$
11:         ${pool}_{v}\leftarrow{pool}_{v}+\gamma_{p}$
12:     else
13:         ${pool}_{v}\leftarrow{pool}_{v}+M_{k}$
14:         $M_{k}\leftarrow 0$
15:         $\mathcal{K}^{t}\leftarrow\mathcal{K}^{t}\setminus\{k\}$
16:for $k\in\mathcal{K}_{m}^{t}$ do
17:     $M_{k}\leftarrow M_{k}+\frac{{pool}_{v}}{|\mathcal{K}_{m}^{t}|}$
18:${pool}_{v}\leftarrow 0$
Algorithm 2 Reward-and-slash design for a set of randomly selected voters.
4.5 Theoretical Analysis on Malicious Votes
In this section, we theoretically show that malicious voters in the proposed framework have no incentive to make dishonest votes.
Theorem 1 (Honest Voting Hypothesis)
When all clients are rational and there is no collusion among malicious clients, a malicious client should not make a malicious vote.
Proof 4.2.
Let $\mathcal{K}_{v}$ denote a randomly selected set of voters and $n_{v}=|\mathcal{K}_{v}|$. For client $k\in\mathcal{K}_{v}$, let $\gamma_{v}>0$ denote the staked tokens for voting, i.e. missing client $k$ must stake $\gamma_{v}$ to participate in the voting, otherwise, it will be removed from the system.
Under A4, each client makes an independent decision on voting. Let $r$ be the ratio of malicious clients in $\mathcal{K}_{v}$, there are $r\cdot n_{v}$ malicious clients in $\mathcal{K}_{v}$ and $(1-r)\cdot n_{v}$ honest clients. If $r\cdot n_{v}<(1-r)\cdot n_{v}$, i.e. missing $r<0.5$, each malicious client will lose $\gamma_{v}$; if $r\cdot n_{v}>(1-r)\cdot n_{v}$, i.e. missing $r>0.5$, each malicious client will gain $\frac{(1-r)\cdot n_{v}\cdot\gamma_{v}}{r\cdot n_{v}}=\frac{1-r}{r}\gamma_{v}$. The expected return $\mathcal{R}$ of a malicious client will be
[…]
Under A2, each client is rational. As $\mathcal{R}<0$, in the long run, a malicious client will lose all tokens and be removed from the system. So, a given client has no reason to make a dishonest vote resulting in honest votes by all clients.
Additionally, there is a game among malicious clients. As Eq. (4) is known by all clients in advance. A malicious client can easily make an honest vote to gain tokens from other malicious clients who make dishonest votes. However, according to Nash Equilibrium [41], we are certain that, in the long run, no malicious clients will make dishonest votes.
Theorem 1 will further be empirically validated in Sec. 5.2.1.
In practice, A4 can be relaxed, where multiple malicious clients work together to attach the FL system. If A2 holds, we are certain that the malicious voters will reach a consensus internally before they act to win the majority vote. Intuitively, all malicious voters can be considered as a group together. In this case, this ”group” will behave exactly as the single malicious client in Theorem 1 based on the same reasoning. The proof is omitted.
4.6 Training
Each round consists of the following steps: proposer selection, local training, global aggregation, local validation, majority voting, token redistribution, and block creation (recording state [^3] For example, the state can record the global model and tokens of each client. information). The above steps are repeated in multiple rounds until certain stopping criteria are fulfilled. The complete training process is depicted in Fig. 3. The stopping criteria could be a fixed amount of training epochs, which is commonly adopted in ML.
5 Experiments
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 4: Token distribution results for malicious voters when all proposers are honest. The malicious voters’ tokens decrease quickly as the number of epochs increases and a large $\gamma_{p}$ leads to a high decreasing rate. This empirically validates our proof of Theorem 1. The solid line denotes the mean over 5 runs with different random seeds and the shaded region denotes 1 standard deviation around the mean.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 5: Federated training under different values of the ratio of malicious clients ($\eta$). The solid lines are the mean AUROCs and the shaded regions are 1 standard deviation around the means. FedAVG w/ Block significantly outperforms FedAVG w/ mal, while being comparable with FedAVG w/o mal, the performance upper bound under this setup.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 6: Global accuracy results when choosing $\gamma_{p}=2,4,8,16,$ and $32$. The global accuracy is not sensitive to the value of $\gamma_{p}$ but the convergence tasks more epochs for larger $\eta$.
Table 1: Performance comparison under different values of the ratio of malicious clients ($\eta$). The reported numbers of the performance are mean and standard deviation under 5 random seeds.
Model $\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.4$
FedAVG w/ mal $0.963\pm 0.017$ $0.946\pm 0.034$ $0.801\pm 0.222$ $0.709\pm 0.266$
FedAVG w/ block (Ours) $0.965\pm 0.008$ $0.969\pm 0.003$ $0.952\pm 0.020$ $0.955\pm 0.021$
FedAVG w/o mal $0.975\pm 0.004$ $0.975\pm 0.004$ $0.975\pm 0.004$ $0.975\pm 0.004$
Oracle $0.971\pm 0.007$ $0.971\pm 0.007$ $0.971\pm 0.007$ $0.971\pm 0.007$
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 7: Token distribution results for clients when setting the parameter for slashing proposers as $\gamma_{p}=8$. The expected average token of malicious proposers fluctuates down during the training process.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 8: Token distribution results for malicious clients when choosing $\gamma_{p}=2,4,8,16,$ and $32$. The expected average token of malicious proposers exhibits a higher rate of decrease when a large value of $\gamma_{p}$ is selected.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 9: Token distribution results for honest clients when choosing $\gamma_{p}=2,4,8,16,$ and $32$. The expected average token of honest proposers displays a higher rate of growth when a large value of $\gamma_{p}$ is selected.
5.1 Experimental Setup
We evaluate the proposed framework in a simulated environment.
5.1.1 Data and Task
We consider a standard classification task, namely loan default prediction. We use the Kaggle Lending Club dataset [^4] https://www.kaggle.com/datasets/wordsforthewise/lending-club to simulate a realistic financial application scenario. We pre-process the raw dataset by dropping all entries with missing values. For the labels, we only keep “Fully Paid” and “Charged Off” to simplify the task as a binary classification task. We randomly select $80\%$ of the data as the training set and use the rest of the data as the test set. The training set is split into $K$ subsets of equal size and distributed across $K$ clients. Within each client, $20\%$ of the local data are randomly selected as the validation set.
5.1.2 Implementation
There are $K=50$ clients in the system and each client is initialized with 64 tokens. We use a 3-layer multi-layer perceptron (MLP) as the network backbone. Apart from the last layer, each layer of the MLP has 128 hidden nodes. We use a standard Adam [42] optimizer with fixed learning rate $10^{-3}$ and batch size 128. No data augmentation is applied. We use the binary accuracy as both the local validation score and evaluation metric. We consider a simple data poisoning attack [14], where malicious clients are trained to confuse the model. All baselines are implemented in PyTorch 1.12.1 [43] on one NVIDIA Tesla T4 GPU. We leverage Ethereum smart contracts to deploy our reward-and-slash design in a private blockchain and simulate the training process using the Python library Web3.py [^5] https://web3py.readthedocs.io/en/v5/. We set $\epsilon=0.05$ based on empirical experience. [^6] We notice that too small $\epsilon$ can cause large oscillation, which slows the convergence, and too large $\epsilon$ can facilitate the convergence at the expense of decreased detection performance, i.e. missing the system fails to remove the majority of malicious clients.
5.1.3 Baselines
We consider 4 baselines. The first one is an Oracle approach, a centralized baseline without malicious attacks. The Oracle should provide the upper-bound performance of the experiment. The second one is FedAVG without malicious attacks (denoted as FedAVG w/o mal), which is equivalent to FedAVG under $\eta=0$ and should provide the upper-bound performance for a decentralized environment. The third one is FedAVG under malicious attacks (denoted as FedAVG w/ mal), where $\eta$ of clients are malicious. The fourth one is the proposed method, FedAVG with blockchain under malicious attacks (denoted as FedAVG w/ block). For FL baselines, $10\%$ of clients are randomly selected to perform local training at each epoch. For FedAVG w/ block, we simply use the remaining $90\%$ of the clients as voters.
5.2 Results
5.2.1 Empirical Analysis on Malicious Voters
To empirically validate the theoretical result in Sec. 4.5, we first simulate a hypothetical scenario where there are only honest proposers. As there are more honest proposers than malicious proposers at each round on average, the effect of malicious weights can be seen as slowing the convergence and decreasing the global performance, which will be validated in Sec. 5.1.3. Note, due to A4, we further simplify the scenario to focus on the behavior of malicious voters. As shown in Fig. 4, given the set of the hyperparameter for slashing voters $\gamma_{r}=\{2,4,8,16,32\}$, the malicious voters will be eliminated from the system shortly (i.e. missing their average tokens decline to 00 within $\approx 40$ epochs).
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 10: Malicious proposers survival time with various $\gamma_{p}$. The expected survival time of malicious proposers declines as $\gamma_{p}$ increases.
(a) $\eta$ = 0.1.
(b) $\eta$ = 0.2.
(c) $\eta$ = 0.3.
(d) $\eta$ = 0.4.
Figure 11: Honest proposers survival time with various $\gamma_{p}$. A large $\gamma_{p}$ can also decrease the expected survival time of honest proposers when the malicious rate $\eta$ is large. This is because, in each epoch, the randomly selected proposers will be all slashed when the performance of the aggregated global model does not increase.
5.2.2 Comparison with Baselines
Following Theorem 1 and Sec. 5.2.1, we now are certain that there will be no de facto malicious voters. Thus, in the following experiments, we focus on the scenarios where malicious clients only upload harmful weights but make honest votes. We evaluate the proposed framework against the baselines described in Sec. 5.1.3. We provide the learning curves in Fig. 5 and the accuracy for all four approaches after convergence (the mean accuracy of the last 50 epochs) in Tab. 1. The performance of FedAVG w/ block is competitive with FedAVG w/o mal (i.e. missing $\eta=0$) and consistently outperforms FedAVG w/ mal. As $\eta$ increases, the performance of FedAVG w/ mal decreases significantly, with a larger standard deviation and increased instability. In contrast, FedAVG w/ block maintains robust performance, with only slightly lower results compared to FedAVG w/o mal.
5.2.3 Analysis of Token Distributions
Fig. 7, 8 and 9 depict the average tokens remaining in honest and malicious proposers during the FL training process. We observe that honest proposers gradually accumulate more tokens while malicious proposers have fewer tokens over sufficient training epochs. Eventually, most malicious proposers lose the ability to participate in staking and are removed from the FL system, as their remaining tokens are insufficient. This meets the expectations of our system design.
5.2.4 Survival Analysis of Clients
As shown in Fig. 10, the anticipated survival time of malicious proposers experiences a decrease as $\gamma_{p}$ increases. This effect can be attributed to the incentive mechanism in place, whereby a higher value of $\gamma_{p}$ results in a greater penalty for proposers who act maliciously. Fig. 11 shows the survival time of honest proposers under different values of $\gamma_{p}$ and exhibits noteworthy behavior. In cases where the malicious ratio $\eta$ is high, the expected survival time of honest proposers may decrease with a large $\gamma_{p}$. This is due to the fact that, in each epoch, all randomly selected proposers will be slashed if the performance of the aggregated global model does not show improvement. Therefore, it is worth noting that balancing the token slashing parameter $\gamma_{p}$ is crucial, because setting an excessively high value can harm honest proposers, whereas a small value can lead to slow convergence (see Fig. 6).
5.2.5 Sensitivity to Malicious Client Ratio
The results presented in Fig. 5 demonstrate the robustness of our proposed method, FedAVG w/ block, against different malicious client ratios, as its performance remains unaffected even under large $\eta$ values. However, it is important to note that the malicious client ratio can impact the token distribution and survival time of clients. Specifically, when there are more malicious clients present in the system, honest clients tend to accumulate more assets on average (c.f. Fig. 9(a) - 9(d)). Nevertheless, they also face a higher risk of being slashed during an epoch, which can ultimately shorten their survival time (c.f. Fig. 11(a) - 11(d)).
5.2.6 Limitations
In this work, as the experimental results aim to evaluate the robustness of the proposed framework, several practical challenges are simplified, e.g. missing staleness [44], storage, and privacy [18]. Further, the proposed method requires more computational power than traditional methods due to mining (blockchain computing) and voting. Finally, large models have gained in popularity in practical applications, e.g. missing ViT [45] and GPT-3 [46]. This raises the question on how to efficiently handle on-chain aggregation for large models. Future work thus will aim to address these limitations to facilitate the research and development of FL with blockchain.
6 Conclusion
In this work, we explore an under-explored research direction, namely using FL and blockchain to defend against malicious behaviors. The defense mechanism is twofold. We use on-chain smart contracts to replace the traditional central server and propose a stake-based majority voting mechanism to detect client-side malicious behaviors. We not only provide a solution to the problem of interest, an emerging direction on trustworthy ML, but also show the robustness of the proposed method and provide the first empirical understanding of the problem.
Acknowledgment
The authors would like to thank Shuhao Zheng from the School of Computer Science, McGill University for the discussion in the early stage.